The mayor of a town has proposed a plan for the construction of a new community. A political study took a sample of 800 voters in the town and found that 52% of the residents favored construction. Using the data, a political strategist wants to test the claim that the percentage of residents who favor construction is more than 48%. Find the value of the test statistic. Round your answer to two decimal places.
Answer #1

Solution :

n = 800

= 0.52

P0 = 0.48

1 - P0 = 1 - 0.48 = 0.52

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

=0.52-0.48 / [0.48*0.52 / 800 ]

= 2.265

Test statistic = z = 2.26